The joint age of uncle Tom and his brother is 60years.In 6years time Uncle Tom will be twice as old as his brother was 6years ago.How old is uncle Tom?(hint :to solve this problem,let uncle Tom be x years old)
Answer by Cromlix(4381)

Hi there,
The joint age of uncle Tom and his brother is 60years
T + B = 60
T = 60 - B
In 6years time Uncle Tom will be twice as old as his
brother was 6years ago.
T + 6 = 2(B - 6)
T + 6 = 2B - 12
Replace T with 60 - B
60 - B + 6 = 2B - 12
Collect like terms
3B = 78
B = 26
Uncle Tom is 34 years old
Brother is 26 years old.
Hope this helps :-)